OpenStudy (anonymous):

Two boys are comparing how fast they are growing. Alex is 60 inches tall and growing at a rate of 1/4 of an inch every month. Leo is 48 inches tall and growing at a rate of 1/2 of an inch every month. Graph the point representing the number of months it will take Alex and Leo to be the same height.

OpenStudy (anonymous):

Here's the graph.

OpenStudy (anonymous):

Are you sure, that they both grow 1 inch each month? Because then they will never reach the same height.

OpenStudy (anonymous):

ohh nvm, it bugged for me. Now i see that it is 1/4 and 1/2.

OpenStudy (anonymous):

I fixed the question, Alex is growing at 1/4 inch a month and lea 1/2 inch a month.

OpenStudy (anonymous):

So you can describe their height as a function. Alexs height: \[f(x)=60+\frac{1}{4}x\] and Leas height: \[f(x)=45+\frac{1}{2}x\]

OpenStudy (anonymous):

Here x is the time in months.

OpenStudy (anonymous):

So now you want to set the 2 equations equal to each other, and then solve x. Then you find how many months it takes for them to be equal height. \[60+\frac{1}{4}x=45+\frac{1}{2}x\] Can you solve x from here?

OpenStudy (anonymous):

Oh, I think I understand it a little more now. Yeah, I'll try it out! If I still have trouble can I tag you?

OpenStudy (anonymous):

Sure, but im going offline in like 10 mins. But sure just tag me

OpenStudy (anonymous):

Okay, thank you so much! :)

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Okay, @Tommynaiter I understand how the equation's supposed to be formatted and what each part means, I'm just having trouble understanding how to solve it haha. Like what are the final steps I take to get the answer?

OpenStudy (anonymous):

Usually I'm pretty good at functions but I'm having trouble with this one.

OpenStudy (anonymous):

So you want to solve x. First subtract 1/4x. This gives you \[60=45+\frac{1}{2}x-\frac{1}{4}x\]\[60=45+\frac{1}{4}x\] Subtract 45 on each side: \[15=\frac{1}{4}x\] Multiply with 4: \[60=x\]

OpenStudy (anonymous):

So this gives you the graph.

OpenStudy (anonymous):

Hope it helped, i gotta go now

OpenStudy (anonymous):

Oh, I was still working on it myself but thank you for the help!